Please help!
Three years ago, pat was half as old as Duane. Seven years from now, Duane will be as many years older than Pat as Pat's age is three years ago. How old is Duane now if he is 4 years older than Pat?

Answer by kapilsinghi(68)

Three years ago, pat was half as old as Duane. Seven years from now, Duane will be as many years older than Pat as Pat's age is three years ago. How old is Duane now if he is 4 years older than Pat?

let present age of Pat = x
so Duane present age = x+4 --------"Duane now if he is 4 years older than Pat"

Ages 3 years ago
-----------------
age of Pat = x-3
age of Duane = x+4-3 = x+1
but "Three years ago, pat was half as old as Duane"
so x-3 = (x+1)/2
2*(x-3) = x+1
2x - 6 = x+1
2x-x = 1+6
x = 7
pats age = 7
Duane age = 7+4 = 11
kapilsinghi123@gmail.com